OpenStudy (anonymous):

use the exponent laws to simplify each of the following until you have only a single power. a) (3^6 x 3^-2)^3 /3^2

OpenStudy (cwrw238):

= (3^4) ^3 / 3^2 = 3^12 / 3^2 = 3 ^ 10

OpenStudy (anonymous):

\[\left(\begin{matrix}2^2 \\ 5\end{matrix}\right)\div \left(\begin{matrix}2^-2 \\ 5\end{matrix}\right)\]

OpenStudy (anonymous):

that is question b) ^